OpenStudy (anonymous):

Can the intermediate-value theorem be used to show there is a solution the equation f (x) = 0 on the interval [0, π⁄2 ]? Give an explanation why. F(x) = 2sin(x) - 3cos(x) -x^2

OpenStudy (anonymous):

a) No. because f (0) > 0 and f ( π⁄2 ) > 0. b) No, because f (0) < 0 and f ( π⁄2 ) < 0. c) Yes, because f (0) > 0 and f ( π⁄2 ) < 0. d) Yes. because f (0) < 0 and f ( π⁄2 ) > 0.

OpenStudy (anonymous):

replace x by 0 and also by \[\frac{\pi}{2}\]

OpenStudy (anonymous):

if one is positive and the other answer is negative, then the function must be zero somewhere in between.

OpenStudy (anonymous):

\[f(0)=2\sin(0)-3\cos(0)-0^2=0-3-0=-3\]

OpenStudy (anonymous):

just taking this all in.

OpenStudy (anonymous):

\[f(\frac{\pi}{2})=2-0-(\frac{\pi}{2})^2\]\]

OpenStudy (anonymous):

the function is continuous yes? so it does not skip values

OpenStudy (anonymous):

if it is say 3 at some point and then -2 at some other point it must take on all values in the interval from -2 to 3. unfortunately in this case both answers are negative

OpenStudy (anonymous):

what is the 2nd value after plugging in pi/2?

OpenStudy (anonymous):

first plug in 0. i got -3 did you?

OpenStudy (anonymous):

then plug in \[\frac{\pi}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got \[2-(\frac{\pi}{2})^2\]

OpenStudy (anonymous):

i got both, since one is pos and the other neg it must cross the x axis

OpenStudy (anonymous):

actually no

OpenStudy (anonymous):

-3 is negative

OpenStudy (anonymous):

I got 2sin(pi/2/0 -3 cos(pi/2) -(pi/x)^2

OpenStudy (anonymous):

and so is \[2-(\frac{\pi}{2})^2\]

OpenStudy (anonymous):

\[F(x) = 2\sin(x) - 3\cos(x) -x^2\] \[F(\frac{\pi}{2}) = 2\sin(\frac{\pi}{2}) - 3\cos(\frac{\pi}{2}) -(\frac{\pi}{2})^2\]

OpenStudy (anonymous):

\[=2-(\frac{\pi}{2})^2\]

OpenStudy (anonymous):

whoops missed the 2, for pi/2 for sin

OpenStudy (anonymous):

got it now!

OpenStudy (anonymous):

but this is negative also. i just checked

OpenStudy (anonymous):

so both values are negative, so we don't know whether it crossed the axis or not

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

answer b is correct.

OpenStudy (anonymous):

therefore IVT cannot explain

OpenStudy (anonymous):

zactly

OpenStudy (anonymous):

got another one, going to do it meow

OpenStudy (anonymous):

thank you, that was very clear!

OpenStudy (anonymous):

i'm ready. let me get my beer

OpenStudy (anonymous):

haha!!!!